Towards the end of a 400m race, Faisal and Edward are leading and are both running at
6m/s. While Faisal is 72m from the finish line Edward is 100m from the finish line.
Realising this and to beat Faisal, Edward decides to accelerate uniformly at 0.2 m/s2
until
the end of the race while Faisal keeps on the same constant speed. Does Edward succeed in
beating Faisal?
1
Expert's answer
2020-12-06T17:18:08-0500
"t=\frac{s_F}{v}=\frac{72}{6}=12\ s"

"s_E=vt+0.5at^2\\s_E=(6)(12)+0.5(0.2)(12)^2=86.4\ m<100\ m"

Edward does not succeed in beating Faisal!